### v1

# Euclid preparation

Euclid preparation
ESA’s Euclid cosmology mission relies on the very sensitive and accurately calibrated spectroscopy channel of the Near-Infrared Spectrometer and Photometer (NISP). With three operational grisms in two wavelength intervals, NISP provides diffraction-limited slitless spectroscopy over a field of 0.57 deg2. A blue grism covers the wavelength range 926–1366 nm at a spectral resolution –900 for a 05 diameter source with a dispersion of 1.24 nm px-1. Two red grisms span 1206 to 1892 nm at –740 and a dispersion of 1.37 nm px-1. We describe the construction of the grisms as well as the ground testing of the flight model of the NISP instrument where these properties were established.
Key Words.:
Cosmology – Euclid – Instrument – NISP – Spectroscopy – optic1 Introduction
Euclid is a new space mission within the European Space Agency’s (ESA) Cosmic Vision 2015–2025 programme, aiming to study the nature of dark matter and dark energy by estimating the expansion history of the Universe and the growth rate of cosmic structures (Laureijs et al., 2011; Euclid Collaboration: Mellier et al., 2025). For this purpose Euclid was designed as a wide-field telescope in the visible to near-infrared (NIR) wavelength range to survey 14 000 deg2 of extragalactic sky from the Sun–Earth Lagrange point 2 (Euclid Collaboration: Scaramella et al., 2022).
To achieve its objectives, Euclid’s payload consists of two scientific instruments developed by the Euclid Consortium: VIS provides high-resolution images of the sky in a single broad wavelength band from 550 nm to 950 nm (Euclid Collaboration: Cropper et al., 2025). With a spatial sampling of and a tight control on the point spread function (PSF), the VIS instrument is optimized for the the measurement of galaxy shapes for weak lensing, which enables the mapping of dark matter structure as tool for cosmological studies (Cropper et al., 2016; Euclid Collaboration: Cropper et al., 2025). Euclid’s second instrument is the Near-Infrared Spectrometer and Photometer (NISP) operating in the NIR (Euclid Collaboration: Jahnke et al., 2025), sharing with VIS a common field-of-view (FoV) of 0.54 deg2 (Euclid Collaboration: Mellier et al., 2025). NISP has been designed for the spectroscopic measurements of galaxy distances for the galaxy clustering probe, as well as NIR information for photometric redshifts.
As a cosmology mission, Euclid’s ultimate goal is to reach a combined cosmological ‘Figure of Merit’ FoM 400 (Euclid Collaboration: Mellier et al., 2025) on the dark energy parameters with weak lensing and galaxy clustering. Reaching this ambitious goal requires a combined approach that leverages both a comprehensive joint analysis of weak gravitational lensing, galaxy-galaxy lensing, and galaxy clustering (commonly referred to as a -point analysis) and a three-dimensional galaxy clustering analysis using spectroscopic redshifts to evaluate cosmological distances. To that end, the Euclid spectroscopic survey is designed to target H line-emitteing galaxies in the redshift range 0.9–1.8, thereby providing precise measurements of the growth of structure and redshift-space distortions as well as the expansion history via baryon acoustic oscillations. Consequently, the galaxy clustering analyses requires to obtain spectroscopic redshift for at least 1700 galaxies per square degree. This corresponds to redshift measurements for 2.55107 galaxies over the survey area of 14 000 deg2 of the Euclid Wide Survey (Euclid Collaboration: Scaramella et al., 2022) achieved through slit-less spectroscopy. The Euclid core science also requires the galaxy redshifts to be estimated with an accuracy , which translates to a minimal spectral resolution when considering a source size of .
In the following we will describe the NISP spectroscopic channel that was designed and built for Euclid’s NIR spectroscopic survey. Section 2 provides a general overview of the NISP instrument as well as the grisms as the key components for NISP’s spectroscopic capabilities. The setup for NISP ground tests of the spectroscopic channel are described in Sect. 3. NISP’s spectroscopic optical quality is presented in Sect. 4 and Sect. 6 shows a detailed analysis of the grisms’ dispersion properties. In Sect. 7 the overall resulting spectroscopic performance is discussed. The paper closes with a brief conclusion and a description of available data products characterising the grisms in Sect. 8.
2 The NISP instrument
The NISP instrument (Maciaszek et al., 2022; Euclid Collaboration: Jahnke et al., 2025) is a near-infrared spectrometer and photometer whose spectroscopic channel has been designed and optimised to detect H emission line galaxies in the redshift range 0.9, 1.8].
The NISP general layout and components are shown in Fig. 1 – a full presentation of the instrument can be found in Euclid Collaboration: Jahnke et al. (2025), and the photometry channel and bandpasses are described in Euclid Collaboration: Schirmer et al. (2022). Here we will provide a brief introduction into the central aspects relevant for spectroscopy.
2.1 Overview of NISP components and properties
NISP consists of a Silicon Carbide (SiC) mechanical structure (NISP Structure Assembly STructure – NI-SA-ST) that houses the detector system (NI-DS) on one side and the optical system on the other (Bougoin et al., 2017). The SiC was chosen for its stiffness and its thermal stability allowing to maintain optical alignement and stability over the mission lifetime. Additionally, fine tuning of internal NISP heaters maintain the opto-mechanics temperature in the required range around – to further maintain optical alignment. Low conductance ( in total) bipods and monopod (NISP Structure Assembly HexaPods – NI-SA-HP) interface the NISP instrument with the the Euclid Payload Module (PLM) baseplate.
The NI-DS (Bonnefoi et al., 2016) acquires images by sampling the FoV with 16 HgCdTe 2k 2k pixels NIR detectors produced by Teledyne Imaging Sensors, arranged in a 4 4 mosaic. Their pixel size of 18 provides NISP with a spatial sampling of the sky of , and the detectors have their wavelength cut-off tuned to 2.3 to minimize thermal noise from the optical system’s thermal emission. Maintaining the focal plane below is crucial for optimal performance and is achieved through thermal coupling of the instrument with an external radiator using four highly conductive thermal straps. Furthermore, NISP’s detectors are read out by 16 individual SIDECAR ASICs, each operating at approximately and dissipating up to of heat. To manage this heat load effectively, high-efficiency thermal coupling to the baseplate is employed, facilitated by methane heat pipes (MHP)s. Due to the substantial heat dissipation, a large radiator of approximately is required to radiate this heat into space, accounting for both the electronics’ heat and any conductive or radiative leaks between the PLM and the radiator.
The NISP instrument images the sky in two different channels: a photometric channel for the acquisition of images with broadband filters and a spectrometric channel for the acquisition of slit-less dispersed images. The optical system of NISP (Bodendorf et al., 2019; Grupp et al., 2012) comprises several elements; a spherical-aspherical meniscus-type corrector lens (CoLA) at the entrance of NISP that actively takes part in correcting the chromatic aberration of the image; the filter wheel (NI-FWA) with a set of three filters (Euclid Collaboration: Schirmer et al., 2022), namely (950–1212 nm), (1168–1567 nm) and (1522–2021 nm) as well as a ‘closed’ position, blocking light from the telescope; the grism wheel (NI-GWA) with a set of three red grisms ( nm), a single blue grism ( nm) and an open position; and finally an assembly of three lenses which, together with their holding structure, constitute the camera lens assembly (CaLA), focussing light onto the detector plane.
2.2 The NISP grisms
The key components in creating dispersed light for determining redshifts of distant galaxies are the NISP grisms. They are the combination of a transmission grating with a prism. This combination allows us to create a common FoV with the NISP photometry channel, placing the dispersed light onto the same detector plane. The NISP instrument contains four different grisms labelled RGS000, RGS180, RGS270, BGS000, three in a red and one in a blue bandpass.
Each of the grisms comprises several optical elements, each fulfilling a specific function (Costille et al., 2019), presented in Fig. 2. A resin-free blazed dispersion grating (dark blue) is directly engraved onto the hypotenuse surface of a fuse-silica prism (light blue) to disperse the light. Each grism has a diameter of 140 mm and a central thickness of approximately 12 mm (Costille et al., 2014), making them the largest grisms ever flown into space. The prism and the grating design were defined to maximize transmission efficiency at the un-deflected wavelength of about 1.2 for the red and 0.9 for the blue grisms, respectively. Additionally, the grating groove has been engraved with a curvature to apply a spectral wavefront correction to the incoming light. The base of the prism carries optical power, that is, it has a curved surface (dashed line) to provide a fine-focus correction for the incoming light beam towards the NISP detection plane. Finally, a multilayer coating (yellow) has been deposited on the curved surface to define each grism’s transmission bandpass.
The grisms RGS000, RGS180 and RGS270 are three red grisms transmitting light in the spectral band , ranging from about nm to nm, with these wavelengths marking half of the maximum of the in-band transmission. The grism BGS000 is a blue grism with a spectral band ranging from about 880 nm to nm. Combined with the transmission characteristics of Euclid’s dichroic and telescope mirrors, the Euclid blue spectroscopic bandpass is reduced to – nm. This work will focus on the grism characteristics themselves, as telescope mirrors and dichroic were not present during the instrument test campaigns. Their – well known – characteristics are simply folded in later.
The grisms create different orientations of spectral dispersion on the focal plane. With a view from the NISP optics towards the projected spectra on the detector plane and with grism RGS000 as the reference, the grism RGS180 has a 180∘ orientation and hence disperses light in the opposite direction as RGS000. Similarly, RGS270 has a 270∘ counter-clockwise orientation with respect to RGS000. These three different orientation were chosen to permit ‘decontamination’ of spectra from overlaps with other sources, creating a reconstructed spectrum from successively acquired exposures of a given field with the three red grisms.
Each grism is glued into a mechanical mount that provides the mechanical interface of the grism with the wheel structure, in order to both provide a defined position as well as minimise stresses from cool-down to the cryogenic environment of the instrument during operations.
3 NISP ground testing setup
After completion of the instrument integration in 2019, the NISP flight model underwent a series of tests in vacuum and with the instrument cooled down to its operational temperature. Two test campaigns were conducted inside the ERIOS vacuum chamber at the ‘Laboratoire d’Astrophysique de Marseille’ (LAM) in the fall of 2019 and the beginning of 2020, to validate the instrument’s performance and its functionality in a space environment.
The ERIOS chamber (Costille et al., 2016b) is a large cryogenic chamber, with an external envelope of 6 m in length and a diameter of 4 m, capable to achieve high vacuum ( 10-6 mbar) and to cool down an inner volume of 45 m3 to low temperatures ( 80 K) with a high stability (4 mK) thanks to liquid nitrogen shrouds covering the entire inner surface of the chamber.
In addition, various types of ground-support equipment were specifically developed by the NISP instrument team to enable the ground tests (Costille et al., 2017). One important component was a point-like light source developed to measure the NISP object plane and to verify NISP’s optical performance, mainly plate scale across the detector plane, point-spread-function width, and a rough estimate of ghost images. The point-like source was the combination of a pin-hole light source and a telescope made of a 160 mm diameter elliptical off-axis mirror with primary and secondary focal points at 500 mm and 3000 mm, respectively, simulating Euclid’s F/20 telescope beam.
The entire aperture of the mirror was illuminated through this 2 pin-hole located at the primary focus of the elliptical mirror, creating a point source for NISP. Both, the mirror and the pin-hole were attached to the same baseplate and the whole system (mirror, pin-hole, and baseplate) was made from silica to ensure stability of the mirror’s focal distance. The telescope baseplate was attached to translation and rotation stages using thermal flexures made from a low conductivity material, thermally isolating the telescope, while providing it with 5 degrees of freedom when being pointed towards different positions in NISP’s FoV.
During the tests, this telescope was inside the ERIOS chamber, together with the instrument, and was operated at temperature of with stability of over a test duration of 1 month. The motors of the telescope mount were kept at warm ambient temperature and were isolated from the cold environment of the ERIOS chamber by a Multi-layer Insulation blanket.
Inside ERIOS, light was fed to the telescope through a cryogenic fibre connected to the pin-hole, with light sources being located outside the chamber. At the interface of the ERIOS chamber a pair of vacuum windows connected the fibre with the light sources, using a cooled neutral density filter to suppress thermal background. On the outside, a set of neutral density filters allowed to select the intensity of light entering the fibre.
Different types of light sources were used during the instrument test campaigns. Among them was a continuum laser source that was either connected to a monochromator, providing a selectable monochromatic wavelength in the range from 0.4 to 2.1 , or to a Fabry–Pérot etalon to emulate an emission-line spectrum with a total of 64 emission lines in the bandpass of the NISP grisms. Alternatively, an argon lamp could be used as a source to provide a well known reference spectrum for comparison with the Fabry–Pérot etalon. The monochromator and the Fabry–Pérot etalon were both characterised before delivery to the ground test equipment.
Spectral lamps were fed into the monochromator to characterise and calibrate it, using HgCd, He, or Cs lamps for the visible band below , and an Ar lamp for the near-infrared domain. The output of the monochromator was scanned using either silicon or germanium photodiodes. The calibration yielded a selectable maximum wavelength accuracy of , and the monochromator bandwidth was measured to be approximately in Full-Width-Half-Maximum (FWHM).
The etalon was re-aligned and re-calibrated in warm conditions with a Perkin Elmer Lambda 900 UV/VIS/NIR spectrometer (‘Lambda 900’) before testing of NISP started. During the tests, thermal sensor monitored the stability of the light sources and the etalon. Figure 3 shows the Fabry–Pérot etalon’s spectrum as it was measured and calibrated with a precision of 0.2 nm using the Lambda 900 spectrometer. A comparison to the grisms’s transmission shows the number of Fabry–Pérot’s peaks available for calibration. The grism channels’ total transmission shown here accounts for the transmission of the CoLA and CaLA lenses, the grisms themselves, as well as the detector quantum efficiency. In this figure neither the Euclid mirrors nor dichroic are accounted for, as both elements were not present in the instrument-level test setup. Note that in Fig. 3, the transmission of the -order components of NISP’s spectrograms, which are below few percent at maximum, have been multiplied by a factor of 10 to make them stand out.
Prior to the start of the optical tests, the focal length of the NISP instrument was measured at cold (NISP focal plane at and NISP optics at ), using a monochromatic light beam with a wavelength of 1000 nm in the photometric passband. In this process, PSF widths were measured from spots created by the point source. This was carried out at five positions in the NISP FoV, where hundreds of individual slightly dithered exposures were taken to over-sample the PSF and to increase statistical accuracy on the PSF width measurement. This was repeated for different distances of the telescope simulator from NISP. During the PSFs acquisitions, the relative distance between NISP and the telescope simulator was accurately monitored with to a laser tracker system that measured the position of ten targets attached to the NISP mechanical structure relative to five targets attached to the back of the telescope mirror (Costille et al., 2016a, 2018).
This set of measurements provided a reference focal plane position for the instrument in the passband. From these measurements an optimal object plane position was identified, using an as-built Zemax model111www.zemax.com (Moore et al., 2004) of the NISP instrument, providing the best image quality in all NISP filter and grism bandpasses. All subsequent instrument tests where then made with the telescope simulator pin-hole positioned on this optimal object plane.
4 NISP spectroscopic optical quality
This section describes the optical quality of the NISP spectroscopic channel, when a grism is moved into the NISP science beam. First, RGS000, RGS180, and BGS000 are discussed, then the special case of RGS270, for which a non-conformance has been identified during instrument tests, rendering it in-operational for science use.
4.1 Optical quality of RGS000, RGS180 & BGS000
The NISP optical quality was assessed by evaluating the radii that encircle 50% and 80% of the PSF’s total energy – in the following these radii will be referred to as EE50 and EE80, respectively. In the spectroscopic channel, this verification was done for every grism of the NI-FWA at three monochromatic wavelengths: 1300 nm, 1500 nm, and 1800 nm for all three red grisms and 900 nm, 1120 nm, and 1340 nm for the blue grism. Additionally, we tested image quality of both RGS000 and RGS180 with a grism wheel rotation offset to validate that offsetting the grism positions by this amount would not degrade image quality. This additional measurement was part of an evaluation of a modified survey strategy to overcome the reduced quality of data from RGS270 (see Sect. 5).
The image quality was evaluated with an analysis of monochromatic PSF images taken at the four corners of the NISP FoV. Because the NISP plate scale of per pixel undersamples the PSF, hundreds of PSF images were acquired around each position, dithering by a tenth of a pixel around the target position in the NISP image plane. This dithering, obtained by moving the telescope on its axes, allowed to spatially over-sample the PSFs and reduced the statistical uncertainties on the EE50 and EE80 measurements.
Before measuring PSF properties, images acquired by the NISP instrument were corrected for pixel quantum efficiency and conversion gain, and a NISP residual thermal background acquired under ‘dark test conditions’, i.e. with the NISP NI-FWA in closed position and all light sources turned off. An automatic procedure, derived from the work of Jiang et al. (2016), was used to locate the PSF within the 2k 2k pixels of the target detectors. Once located, a 20 20 pixel stamp centred on the expected PSF position was extracted from the images and PSF metrics were extracted by image analyses. Every extracted sub-image was controlled by eye and e.g. detector cosmetic defects misidentified as PSFs by our algorithm were rejected from the analyses. An Intra Pixel Capacitance (IPC) correction, relying on a Richardson–Lucy deconvolution method (Lucy, 1974; Richardson, 1972) using IPC coefficients previously measured during detector characterization as the deconvolution kernel, was applied to each extracted stamp. However, it should be noted that the spread of the measured EE50 and EE80 distributions were much wider than the correction factor applied through this IPC deconvolution. The IPC signal leaking to neighbouring pixels was measured to be very small, of 2.870.01% losses in the central pixel that are non-equally redistributed to the eight immediately adjacent pixels (Le Graët et al., 2022).
Figure 4 shows one example of a NISP PSF acquisition. The left panel is a full frame image (i.e. 2k 2k pixels) from one detector, and the artificial point source at the top. A ring-like structure is visible surrounding the PSF, that resulted from scattered light at the exit of the pin-hole. Since the ring radius is very large compared to the PSF diameter, it was easily disregarded in the analyses. However, these rings were quite useful in validating that the blind search algorithm indeed successfully located the PSF. In the centre panel of Fig. 4 a zoom into the small box around the PSF is shown. The diffuse circular feature to the right of the compact PSF is produced by light emission coming from the fibre core passing through a non-perfectly blocking neutral density filter around the pin-hole. This was meant to block light throughout the full NISP wavelength range. It however appeared that instead it had a transmission increasing with wavelength, starting with an optical density 5 at 900 nm and reaching an optical density of 4.5 at 1800 nm, creating this patch. The rightmost panel shows another zoom step into the PSF, with contrast adjusted to prevent display saturation.
The EE50 and EE80 are deduced from PSF model fitting with the following model
where are the spatial coordinates in the pixel mosaic reference frame, is a PSF model of parameters – for the functions that were considered see below – is a -Bell function (Boyd, 2006) of parameters – which is used to model the contribution from the fibre core – and , a constant introduced to account for a residual constant background. As the fit is limited to a small stamp of 20 20 pixels centred on the PSF, a constant background is a good first approximation at these scales. Note that the axis is defined such that the light dispersed by RGS000 along a spectral trace has its wavelength increase with .
To estimate the reliability of our estimates, we tested three different functions to model the PSF profile.
An asymmetric-Gaussian function for which amplitude, centroid position, width along the and axes as well as inclination with respect to the axis are free parameters. When using this model, EE50 is evaluated as , with the width of the PSF in the cross-dispersion direction. We used the cross-dispersion direction to minimise the impact of the monochromator bandwidth on the estimate of EE50. However, such an estimation might be underestimating the true width, by assuming the PSF is having a Gaussian profile and neglecting potential contributions from a non-Gaussian tail.
The sum of two asymmetric-Gaussian functions, hereafter refereed as a dual-Gaussian profile, for which the respective amplitude, common centroid position, individual width along the and axes, as well as common inclination are free parameters. To avoid any degeneracy in the model fitting, the amplitude and width of the larger Gaussian was defined by a multiplicative factor relative to that of the narrower Gaussian. This model tends to better account for potential non-Gaussian tails in the PSF profile.
Because the asymmetric-Gaussian profile may not be able to catch the faint extended tail of the PSF, we did not use an analytical estimate of EE80, as we used for the EE50, but instead evaluated the EE80 after having subtracted the contributions from the fiber-core and background in the image, both derived from the model fitting. In this case EE80 is estimated by searching for the radius of the circle – centred onto the PSF centroid position – that encapsulates 80% of the total signal in the background-free image. However, when working with the two other models, to avoid introducing errors when subtracting the background or fibre-core model from the PSF image, the EE50 and EE80 were instead evaluated on the fitted model. In those cases, the radii were determined by evaluating the apertures containing 50% and 80% of the total signal from the PSF model, using either the dual-Gaussian or Moffat components of the model, and excluding both background and fibre-core contributions. However, as the monochromator did not have infinitely narrow bandwidth, and hence PSFs are slightly extended along the dispersion direction, both the EE50 and EE80 are slightly overestimated as this estimation disregards PSF asymmetry.
The reduced -distributions of the three models have a median value of 1.04 for the asymmetric-Gaussian model, 0.97 for the dual-Gaussian model, and 1.02 for the Moffat profile. These values are too similar to prefer one model over the others. For every model a few cases show values larger than 1000. These were found to be correlated between the three models and are associated with the PSF being located close to hot or pixels with ‘defects’ like cosmic ray hits, Random Telegraph Noise, etc… that were not properly accounted for in our data reduction. We subsequently excluded these cases with biased fit results, removing PSFs with reduced which result in excluding about 5% of the PSFs. Additionally, we rejected any PSF with an unphysically large maximum signal 10 000 e- s – given the photon flux provided by the optical ground equipment setup.
Figure 5 shows the relative differences in the measured position of the PSF between the different PSF models. With a standard deviation lower than 0.1 pixels and a mean of (7–10) 10-3 pixels, depending on the histogram, we concluded that all three models predict the same PSF centroid position with an error on the order of one tenth of a pixel. As those three models locate PSF at roughly the same position, it is safe to consider they all identify the same PSF which allowes a comparison of the three profiles.
Despite the differences in the profile definition, the three models provide similar EE values. Given the goal to assess the NISP spectroscopy image quality from PSF size we conclude that these models overall provide robust estimates.
In Fig. 6 the overall image quality is shown, expressed as mean and spread of PSF EE50 and EE80 across the NISP FoV, derived from the three PSF models. These are compared to an ideal optical diffraction-limited system (shaded area) and to the scientific requirements (dashed area). The boxes for each model represent mean, and first and third quartile of their values across the FoV – they are the combined result of the accuracy of our measurement and the actual PSF-size variations. Similarly, the width of the ideal model band shows the variations of the theoretical PSF across the FoV. One can see from this figure that the image quality of the NISP grisms is close to the theoretical expectation from an ideal system, with EE50 0.6 times smaller than the scientific requirements at 1500 nm. A more detailed investigation of the four target fields shows variations in the median value of both EE50 and EE80, suggesting a smooth variation of PSF size across the FoV. However, these are smaller than the spread due to measurement uncertainties and are being disregarded.
To make sure that NISP grisms could also be operated with a rotational positioning offset of without an impact on image quality (for the rationale see next section), additional acquisition with tilted grisms were made during the ground test campaign. Figure 7 compares the image quality for RGS000 and RGS180 for their nominal position and offset. There are no differences. For clarity, this figure is only using the asymmetric-Gaussian profile as it has the smallest dispersion in the estimated EE50 – but since all models lead to similar estimates of EE50, this result also holds for the other models.
One has to keep in mind that the data presented here were taken without the Euclid telescope system, and hence the telescope PSF has to be accounted for when extrapolating the measured NISP performance to in-flight conditions. Ground test of the Euclid payload module, which contains both instruments as well as the whole Euclid optical system, were conducted in summer 2022 at ‘Centre Spatial de Liège’ (CSL) by Airbus Defense and Space. Although these tests are outside the scope of this paper, we can report that while Euclid’s telescope was prone to gravity which stressed the mechanical structure of the primary mirror, analysis did not revealed any degradation beyond expectations.
5 Addressing RGS270 Non-Conformity
One consequence of the Euclid telescope (Refregier et al., 2010; Laureijs et al., 2011; Euclid Collaboration: Mellier et al., 2025) having an off-axis three-mirror Korsch design (Korsch, 1977) is that in order to obtain high-quality images at all positions the NISP focal plane has to be tilted by an angle of with respect to the optical axis (Euclid Collaboration: Jahnke et al., 2025), as illustrated in Fig. 8. The grisms RGS000 and RGS180 are dispersing light in the direction parallel to the axis, i.e. perpendicular to the focal plane tilt. For this reason the focal length for exactly (and only) these two specific dispersions directly is identical for all wavelengths of a given object, as in an on-axis telescope. On the other hand for grism RGS270, with a dispersion direction perpendicular to RSG000 and RSG180, the dispersion is running in a direction perpendicular to the axis, i.e. in the direction of the focal plane tilt. The curvature of the grating grooves in the RGS270 grism design was introduced specifically to correct for the tilt of the NISP focal plane, creating in-focus images on the focal plane at any wavelength.
However, during our ground test campaign at instrument level we observed that images from grism RGS270 were partially defocused. This can be seen in Fig. 9 which compares Fabry–Pérot spectrograms acquired with RGS000 and RGS270 during NISP ground test campaign. Focusing on the middle image presenting the RGS270 2D spectrogram, one can see that the bluest part of the spectrum (left-hand side) has a narrow and well focused PSF. In contrast to this, with increasing wavelength towards the right, the PSF gets successively wider and at the reddest end even shows a ring-like appearance. There is clearly an increasing defocus with increasing wavelength.
After investigations, we concluded that the RGS270 grism was built using an improper interpretation of its optical design: the definition of the RGS270 optical reference frame was not properly propagated to the NISP mechanical reference frame – which are oriented inversely to each other. The main consequence of this error resulted in all models built for RGS270 to disperse light in the opposite direction to the design definition. As a result the focus compensation by the curved grating groove was going in the wrong direction.
A ‘Tiger Team’, composed of Euclid Consortium members, investigated potential hardware solutions, including various possibilities to replace the RGS270 grisms that were built, as well as options to modify Euclid’s surveys. After an assessment of both the risks for the mission induced by dismantling the NISP instrument as well as the time required to build a new RGS270 grism, a survey solution was proposed by the consortium and approved by the ESA. This was based on a detailed analysis of a modification of survey and observation parameters, which we will discuss in the following.
To obtain the additional spectral orientation angles needed for decontamination of overlapping spectra, the survey solution involved to modify the survey strategy using NISP settings that rotate the NI-GWA by away from its nominal positioning of the grisms with respect to the centre of the beam, when observing either with RGS000 or RGS180. This in turn creates spectra tilted by vs. the nominal orientation of spectra from RGS000 and RGS180. After an in-depth simulation of different survey strategies we ended up with an optimal spectroscopic observation sequence. Each field is observed four times, with grisms used in the following order: RGS000 RGS180+ RGS000– RGS180.
The advantage of this observation sequence, executed in the ‘K-pattern’ of telescope dithering offsets for each sky position (Euclid Collaboration: Scaramella et al., 2022; Euclid Collaboration: Mellier et al., 2025), creates a geometry formed by the stacked spectra that offers even more dispersion angles than the initial observation sequence, for disentangling spectra during the offline data processing on ground. The largest drawback is that by slight rotation angle of the grism wheel, the nominal aperture of the grism – defined by a baffle installed in front of the grating – is shifted away from the nominal beam centre, introducing slight vignetting at the edge of the FoV. This vignetting has been estimated from Zemax simulation to result in 10% flux losses at field angle, i.e. distance from the centre of the FoV, which is reduced to flux losses at field angle . On-sky simulation, involving the official Euclid data reduction pipeline, demonstrated that this vignetting level is largely compensated by the detector quantum efficiency and an optical throughput which are both higher than scientifically required, and that the spatial structure can also easily be accounted for in the pipeline. Larger tilts were also investigated but the increasing vignetting level quickly reduces the options for modifying the Euclid survey strategy along this line.
As a conclusion, the RGS270 is kept on board the NISP instrument to maintain the centre of gravity during mission operation, but it will not be used for Euclid’s surveys. While an extensive dataset was acquired for RGS270 during ground testing, in the following the performance of RGS270 will not be discussed any further.
6 NISP spectroscopic dispersion
An important part of the NISP ground test campaign was to establish a preliminary assessment of the grism’s dispersion law as mounted inside the instrument, in preparation for the instrument calibration in-flight and NISP’s in-flight operations. This section is detailing the procedure used to calibrate the NISP spectral dispersion on ground, explaining how the calibration datasets were acquired and analysed, before we discuss the validity and precision of this calibration.
6.1 On-ground spectral dispersion calibration
To estimate spectral dispersion functions of the NISP grisms, the Fabry–Pérot etalon spectrum was projected with the telescope simulator to 144 individual points on the NISP FoV. For every pointing a short exposure of 13 s was taken before the telescope simulator moved to the next field point. Additionally, at all positions corresponding to the centre of a NISP detector, we also used the Argon spectral lamp to obtain spectra as reference and control.
To avoid any bias induced by a small but still different alignment of the grism wheel after multiple repositioning, all measurements at these 144 points were made without turning the grism wheel. Only once a 144-pointing scan with a given grism was completed, the grism wheel was moved to position the next grism. This test sequence was used for all four grisms – meaning the grisms BGS000, RGS270, RGS000, and RGS180 each at nominal positioning as well with the grisms RGS000 and RGS180 with a rotation offset. Also, from the point of view of the calibration, we chose to consider the rotated grism positions RGS000 and RGS180 to be independent grisms of their own and not to consider them to be identical to RGS000 or RGS180.
Before going into the details of the calibration, Fig. 10 present examples of the -order spectrogram for the blue and red grisms acquired during the NISP ground tests. The middle panels each show and compare spectrograms of the Fabry–Pérot etalon with those of the Argon spectral lamp. The top panels show reference Argon spectra in vacuum in these wavelength ranges, taken from the NIST database (Kramida et al., 2022). By comparing the reference Argon spectra with those observed with NISP (middle and bottom panels) one can identify the strongest Argon emission lines present in both the and bandpasses. The middle panels also identify the Fabry–Pérot transmission peaks. For clarity, we limit the labelling to two peaks but based on the Fabry–Pérot calibration data all other transmission peaks were clearly identified222For reference see the calibrated Fabry–Pérot Spectral Energy Distribution (SED) in Fig. 2. The middle and bottom panels do not cover the spectrogram of the -order which is located about 500 pixel away off the left-hand side.
One note, upon close inspection of the bottom panels, that both show a slight periodic excess. This is a charge persistence signal from the previous Fabry–Pérot etalon exposure. However, persistence did not affect the exposure of the Fabry–Pérot etalon itself, since the pointing position was changed between subsequent exposures and the Fabry–Pérot light source was brighter than the spectral lamp.
One central calibration for the NISP grisms is to describe the relation to convert wavelength for a source at any position into a position onto the focal plane. This is first established independently for each of the 144 spectrograms by modelling the 2d spectral trace with the following parametric function:
Here are spatial coordinates of a spectral feature on the focal plane, with and axis defined to be parallel to RGS000’s cross-dispersion and dispersion direction, respectively. are spatial reference coordinates, , are Chebyshev polynomials of the first kind defined in the wavelength range, and the Chebyshev coefficients, with being either or .
The coordinates of the Fabry–Pérot transmission peaks were measured by fitting the observed peaks with the PSF profile described in Sect. 4, providing position with an accuracy of of a pixel.
This work uses as a reference the centroid position of the spectral -order which is defined to be the position of the in-band -order’s wavelength with minimal transmission. This reference position is estimated from a template fit of the 2D image of the -order spectrogram with a Fabry–Pérot template.
The transmission of the optical ground equipment has some residual knowledge uncertainties. To limit the impact of this, the Fabry–Pérot template is constructed by taking the averaged profile of its -order spectrogram measured with the corresponding grism, corrected by the -order transmission efficiency. In this way the template is considered to be representative of the photon flux at the entrance pupil of the NISP instrument. To account for the -order, the template is then multiplied by the -order transmission efficiency (see Fig. 3) and convolved with a PSF profile before being adjusted to the 2D images of the -order. Parameters of the fit are the centroid position of the template, defined to be the central wavelength of the grism band-pass, where the -order transmission efficiency is at its minimum, a stretch factor converting wavelength to pixel, the dispersion direction angle, and parameters of the PSF profile defined in Eq. 1 which are accounting for the contribution of the fibre core.
An example of such a template fit is shown in Fig. 11 which presents the measured image of the Fabry–Pérot -order spectrogram in (A) with the model in (B) and the residual in (C) and (D). Due to the prisms onto the NISP’s grisms, the -order are slightly dispersed, extending over several tens of pixels. Additionally, the grating’s blaze function is optimized to maximize transmission in the -order – around for the red grisms and for the blue grism – while minimizing transmission at the same wavelengths for the -order. As a result, instead of appearing as a point-like feature, as would be expected for a purely dispersive grating, NISP’s -order takes the form of an elongated, double-peaked structure. We tested the different PSF profiles we listed in Sect. 4 and concluded that they all provide similar results without any of them outperforming the others. As for modelling the monochromatic PSF, the usage of either of the profiles described in Sect. 4 led to similar centroiding of the -order with negligible impact on the derived calibration parameters. The -order position errors at were determined using -profiling and subsequently propagated to the centroid position of the Fabry–Pérot transmission peaks during the calibration of individual -order spectrograms. Analyzing the distribution of measured uncertainties in the -order positions, we obtain mean uncertainties of with a standard deviation of pixels in the cross-dispersion direction and with a standard deviation of pixels in the dispersion direction. This highlights that positioning is more precise in the cross-dispersion direction, as the profile of the -order is significantly thinner in this direction, leading to better constraint.
Coefficients of the Chebyshev polynomial as described in Eq. 2 are evaluated from a recursive -fit. In the recursion, outliers were rejected at every iteration, each identified as spectral features located 5 away from the fitted spectral trace. Recursion stopped when no further outlier were identified. This allowed us to automatically reject some of the Fabry–Pérot transmission peaks that were insufficiently characterized by our PSF modelling due to nearby hot or bad pixels. Chebyshev polynomials where expanded up to the third order in the cross-dispersion direction and up to the fourth order in the dispersion direction. For the blue grism calibration the expansion was limited to the third order in both directions. The level of the expansion order was chosen in order for the averaged reduced to be the closest to . With the above expansion order, the averaged reduced is of the order of . However, we noticed that the distribution is skewed toward lower values. This is due to the limited accuracy of the template fit on a few of the -order images which propagated to the -order spectra and leads to large uncertainties in the relative position errors of some peaks’ positions, .
Up to this point, the Fabry–Pérot spectrograms were modelled independently of each other, leading to calibration coefficients that smoothly vary across the FoV depending on the location of the order. To complete the calibration, the spatial dependency of each of the coefficients is then fitted by a bi-dimensional Chebyshev polynomial defined as
with the calibration parameters to be evaluated and the Chebyshev polynomials of the first kind, with being either or defined in the range as
Here are the limits of the focal plane array, set to for both the and axes. Again, the evaluation of the parameters is carried out by a fit. The tables that summarise the calibration coefficients for each of the grisms are presented in Appendix A.
Theoretically, in a perfect optical system with constant grating groove, one expects the spectrogram to rotate around their un-deflected wavelength when the grating is rotated around its centre. However, with the complexity of the NISP instrument itself, the mechanical uncertainties on the wheel position after its rotation, and on the pointing of the telescope simulators, it was not possible to precisely measure and confirm the position of the rotation centre for each of the spectrograms. This choice was made to minimize the bias induced by the uncertainty on the location of each spectrum’s centre of rotation. We tested the hypothesis that the rotation centre of the spectrograms was either their undeflected wavelength or the -order333Which, strictly speaking, is not physically possible given the grism design, but this was nevertheless also tested, since at first approximation the -order could be a proxy for the rotation centre.. However, the calibration error derived from such a nominal position was found to be outside of our scientific requirement once it was applied to the rotated grisms, despite being calculated at sub-pixel level.
Once we obtained the calibration parameters, we computed a predicted position of the Fabry–Pérot emission line using Eqs. 3 and 2 and compared the reconstructed and measured positions. This comparison is shown in Fig. 12. For clarity, we only show the results of the comparison for the blue grism BGS000 (in blue) and for the combined distribution of RGS000 and RGS0004 (in yellow) as well as for the combined distribution of RGS180 and RGS1804 (in purple). The calibration error we derived from this comparison is found to be lower than (i.e. pixel) in the cross-dispersion direction and lower than (i.e. pixel) in the dispersion direction for each grism.
Additionally, we noticed that every spectrum for which the and orders were located on two different detectors shows a local offset in the estimated values of both the and coefficients, while the higher order coefficients of the parametrisation remained compatible with the coefficients estimated for any spectrogram fully falling onto a single detector. This offset behaves as if those spectra have a longer separation between the position of their and orders, breaking the otherwise rather smooth variation of the and coefficients across the field of view. Instead, we concluded that this offset has its origin in a bias induced by a limited precision of the detector metrology: this provides physical positions of the NISP detectors within the focal plane and was used to convert position measured in pixels on a given detector to the position in the NISP focal plane reference frame. However, the detector metrology was only determined at room temperature by measuring the position of reference marks engraved on the mechanical structure of each detector, using an optical camera mounted on metrology bench. This method only provided a precision of the order of one pixel (i.e. ). The NISP focal plane metrology at operational temperature was therefore computed relying on the thermal model of the NISP instrument to predict, with limited precision, how thermal stress would modify the focal plane.
We did try to re-calibrate the detector metrology at cold using the spectroscopic data themself by aligning -order spectra dispersed over multiple detectors. However, the small number of sources available (i.e. one single spectrum per NISP exposure in the entire NISP focal plane) did not allow a re-computation with a sufficient precision for the spectroscopic calibration. We therefore left our cold reference-metrology as is and checked that the calibration parameters remain within 3 when computing calibration parameters with or without the spectra showing an offset in the and coefficients.
6.2 Validation of the spectroscopic dispersion law
Grism | Cross-dispersion | Dispersion | ||||
Average [px] | Std-Dev [px] | Bias error [px] | Average [px] | Std-Dev [px] | Bias error [px] | |
BGS000 | 0.29 | 0.33 | 0.36 | 0.53 | ||
RGS000 | 0.12 | 0.13 | 0.21 | 0.29 | ||
RGS180 | 0.14 | 0.15 | 0.33 | 0.38 |
During the NISP ground test campaign, we acquired 16 spectrograms of the Argon spectral lamp with each of the BGS000, RGS000, and RGS180. An Argon spectrogram was observed every time the telescope simulator pointed to the centre of one of the 16 NISP detectors. This was achieved by replacing the Fabry–Pérot’s fibre feeds within our optical ground equipment with feeds connected to the Argon spectral lamp. No other modifications were made to the instrument configuration or telescope simulator pointing. These Argon spectrograms were used to validate the spectroscopic dispersion law that we derived from the Fabry–Pérot etalon calibration.
Similarly to what was done for the Fabry–Pérot’s spectrogram, we extracted Argon emission line positions based on model fitting of the 2D emission line profile. However, the detection process was significantly impacted by a substantial level of detector persistence charge signal from the previous Fabry–Pérot’s exposure acquired at the same pointing as for the spectral lamp acquisition. A visual inspection of the initially identified line was conducted, to reject persistence signal peaks that are miss-identified as Argon emission lines by the automated peak-search algorithm, as well as ambiguous line detection. In the cases where it was challenging to differentiate between actual emission lines and persistence signal, the line was masked from the analysis.
Additionally, we extracted the -order position from a template fit, replacing the Fabry–Pérot template with the Argon SED template whenever the -order was visible on the 2D spectrogram. This was primarily possible for acquisitions with the BGS000 grism. Contrary to this, the signal was below the detectable signal-to-noise ratio for the red grisms. Since we did not move the telescope simulator when changing the light source, we assumed for those cases that the -order position of the Argon spectrogram was the same as that of the Fabry–Pérot spectrogram from the previous exposure.
When inspecting Argon spectrograms captured with the BGS000, we found that four of them had very weak signal for the -order resulting in the template fit failing to converge, positioning the -order at about seven pixels away from visually determined position. These spectra were also excluded from our analyses.
Using calibration parameters that were derived from the Fabry–Pérot spectrogram analysis (see Sect. 6.1), we estimated the position of the Argon emission lines relative to the -order position and compared the predicted positions with the measured position.
The bias resulting from this comparison is reported in Table 1, defined as the average of the distribution as shown in Fig. 13. It is induced by the persistence present in the images that tend to either bias the centroiding of the Argon lines themselves or the estimate of the position of the -order. As was described above, for the BGS000 an Argon SED template was used to evaluate the -order position, while for both red grisms this was taken from the previous exposure, assuming the telescope did not move between exposures. As a consequence, the weakness of the -order and the persistence signal present in the BGS000 spectrogram leads to a larger bias than what is observed for the corresponding distribution of the red grisms. Nonetheless, despite this bias reported in the Table 1, we found that the predicted position of the Argon lines fall within a distance of pixels of the measured lines, which validates our calibration parametrisation.
7 NISP Spectroscopic performance
While calibration parameters have been computed on the basis of Chebyshev polynomials, it is more convenient to convert the obtained spectroscopic dispersion laws into regular polynomials to assess properties and performance of the grism’s dispersion. In that case, the dispersion laws take the following form
where , with being either or , are the regular polynomial coefficients in cross-dispersion () and dispersion direction (). When converted to regular polynomials, the coefficients of zero degree () are related to the Cartesian distance that separate the -order of the spectra to the reference used in this work, i.e. the position of the -order. The first degree coefficients () are related to the dispersion of the spectra while coefficients of higher degree define the curvature of the spectra, which is linked to the non-linearity of the dispersion law and optical distortions of the system.
For clarity and simplicity, the forthcoming discussion will be confined to the description of the dispersion law averaged over the NISP FoV. Table 2 reports the dispersion coefficients in both the Chebyshev and regular basis and reveals that when the grisms are centred onto their nominal position, the cross-dispersion coefficient is about two orders of magnitude lower than the dispersion coefficient . This aligns with the intended design of the grisms which primarily focuses on dispersing light in a single direction. Optical and mechanical distortions as well as the manufacturing limitations contribute to the distortion of the spectral track in both directions. However, even with those limitations, the higher-order coefficients ( with ) are small enough for the dispersion to be nearly a linear dispersion law mostly directed towards the axis.
In contrast, when the grisms are tilted by , the cross-dispersion coefficient increases to be one order of magnitude lower than the dispersion coefficient . This discrepancy with the nominal position arises due to the dispersion occurring in both directions when the dispersion grating is rotated by .
Under the formalism used in this work, the dispersion of NISP’s grisms predominantly takes place in the direction. As a consequence, our focus will now shift to the discussion of the regular polynomials.
In the context of slit-less spectroscopy, spectral resolution is defined by a combination of both the intrinsic instrument PSF as well as the characteristics of the targeted object’s size, shape, and surface brightness distribution. Referring to the definition provided in Euclid’s scientific requirement, we define NISP’s resolving power as
In the above equation, corresponds to the per-pixel spectral dispersion along the dispersion direction444along the -axis for RGS000 and RGS180 in nominal position and in a mixture of and axis when those grisms are at inclination and is expressed in /px. The corresponds to the full-width-half-maximum of the target source with an axisymmetric Gaussian profile of effectives variance . This effective full-width-half-maximum results from the convolution between the luminosity profile of the source and the PSF and can be expressed as:
The is the full-width-half-maximum of the PSF, in pixels, at a given wavelength , and corresponds to the full-width-half-maximum, in pixels, of the target and depends on the source’s surface brightness distribution.
When considering grisms RGS000 and RGS180 in their nominal position, the regular polynomial given in Eq. 5 describes the parametrisation for the grism’s dispersion law in the dispersion direction. Its first-order coefficient can be related to the grism’s dispersion . To obtain an estimate of the per-pixel dispersion one needs to first invert this polynomial. Given our parametrisation, is continuously differentiable and non-null in the bandpass of the grisms, hence is invertible in the neighbourhood of , with its inverse being also differentiable at such that
The constant term in the derivative of the inverse function gives an estimate of the dispersion at a wavelength within in the grism’s bandpass.
Figure 14 presents the variation across the field-of-view of the dispersion coefficients at the mid-bandpass wavelength, for all combinations of grism/rotation positions that were tested. The dispersion smoothly varies across the NISP FoV following a radial gradient with its maximum, or minimum, on the left-hand side of the focal plane at about mosaic coordinates . Notice the difference between the two red grisms: RGS000 has an average dispersion of the order of while RGS180’s average dispersion is of the order of . The negative sign of the RGS180 dispersion originates from its design which imposes a dispersion of the light in the opposite direction from RGS000. But on overall, RGS000 and RGS180 have a very similar absolute dispersion in term of value and profile over the focal plane array, leading to a similar performance for both grisms. Note that, to highlight these similarities and due to the negative dispersion value of RGS180, color scale for RGS180 is reversed relative to that of RGS000. In contrast, BGS000 has a lower dispersion with an average value of the order of . All these values are compatible with the design-specifications for each grism.
Combining this estimate of the dispersion with the measured PSF size, NISP’s resolving power can be derived using Eqs. 6 and 7. The result is summarized in Fig. 15 which shows NISP’s averaged resolving power for a source of 05 FWHM. The averaged resolving power from the blue grism increases from about at to at , while both red grisms see their resolving power increase from at to at . These values have to be compared to the NISP scientific requirements of a resolving power for the blue grism and for the red grisms and for a typical source size of . The combination of the high optical quality of the grisms and the NISP camera and collimator lenses ensures that NISP fulfils its requirements in term of spectral dispersion and resolving power.
Again, shorter test were conducted at the payload module level with the addition of the Euclid mirrors and dichroic element to the light path. This allowed us to verify that the addition of the folding mirrors as well as the transmission of the near-infrared light beam through the dichroic did not significantly modify NISP’s spectral dispersion and resolving power.
8 Conclusions and available data products
The NISP instrument integration was completed in 2019. Before its delivery to ESA for its integration into the payload module of the Euclid spacecraft, the NISP instrument underwent a series of tests in a vacuum and cold environment to validate its operation mode as well as to validate its performance in view of the forthcoming flight operation. During the test campaigns that took place in the beginning of 2020 the optical performance of NISP were verified.
One of the measurement campaigns conducted under vacuum at cold temperatures was dedicated to the verification of the image quality provided by NISP’s grisms as well as to the verification and validation of their dispersion. These are key properties in the accurate determination of cosmological redshifts which sits at the core of the derivation of cosmological parameters in galaxy clustering analyses.
While a non-conformity was revealed for the grism RGS270, which resulted in the modification of the Euclid survey strategy, the analysis of the ground dataset presented in this work shows that NISP’s grisms BGS000, RGS000, and RGS180 have an extremely high image quality with a PSF encircled energy size well below the scientific requirements and very close to that of an ideal, solely diffraction-limited instrument.
Secondly, these tests allowed us to verify that the spectral dispersion of NISP lies above requirements and showed that NISP’s resolving power is larger than from to . The RGS000, RGS180, RGS270, and BGS000 throughput for the and spectral orders at the centre of the NISP focal plane array are available at an ESA server555https://euclid.esac.esa.int/msp/refdata/nisp/NISP-SPECTRO-PASSBANDS-V1. Available throughput numbers include contributions from the grisms, the telescope, and NI-OA, as well as the mean detector quantum efficiency. For convenience, the tables are available in both ASCII and FITS format.
Acknowledgements.
The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org).References
- Bodendorf et al. (2019) Bodendorf, C., Geis, N., Grupp, F., et al. 2019, in Astronomical optics: Design, manufacture, and test of space and ground systems II, ed. T. B. Hull, D. W. Kim, & P. Hallibert, Vol. 11116 (SPIE), 111160Y
- Bonnefoi et al. (2016) Bonnefoi, A., Bon, W., Niclas, M., et al. 2016, in Space telescopes and instrumentation 2016: Optical, infrared, and millimeter wave, ed. H. A. MacEwen, G. G. Fazio, M. Lystrup, N. Batalha, N. Siegler, & E. C. Tong, Vol. 9904 (SPIE), 99042V
- Bougoin et al. (2017) Bougoin, M., Lavenac, J., Pamplona, T., et al. 2017, in International Conference on Space Optics — ICSO 2016, Vol. 10562 (SPIE), 1329–1337
- Boyd (2006) Boyd, J. P. 2006, Journal of Scientific Computing, 29, 1
- Costille et al. (2016a) Costille, A., Beaumont, F., Prieto, E., Carle, M., & Fabron, C. 2016a, in Advances in Optical and Mechanical Technologies for Telescopes and Instrumentation II, Vol. 9912 (SPIE), 1390–1407
- Costille et al. (2018) Costille, A., Beaumont, F., Prieto, E., et al. 2018, in Advances in Optical and Mechanical Technologies for Telescopes and Instrumentation III, Vol. 10706 (SPIE), 498–511
- Costille et al. (2014) Costille, A., Caillat, A., Grange, R., Pascal, S., & Rossin, C. 2014, in Advances in Optical and Mechanical Technologies for Telescopes and Instrumentation, ed. R. Navarro, C. R. Cunningham, & A. A. Barto, Vol. 9151, International Society for Optics and Photonics (SPIE), 91515E
- Costille et al. (2019) Costille, A., Caillat, A., Rossin, C., et al. 2019, in International Conference on Space Optics — ICSO 2018, Vol. 11180, 1118013, series Title: Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series
- Costille et al. (2017) Costille, A., Carle, A., Beaumont, F., et al. 2017, in International Conference on Space Optics — ICSO 2016, Vol. 10562 (SPIE), 937–945
- Costille et al. (2016b) Costille, A., Carle, M., Fabron, C., et al. 2016b, in Proceedings of SPIE - The International Society for Optical Engineering, Vol. 9904, iSSN: 1996756X
- Cropper et al. (2016) Cropper, M., Pottinger, S., Niemi, S., et al. 2016, in Space Telescopes and Instrumentation 2016: Optical, Infrared, and Millimeter Wave
- Euclid Collaboration: Cropper et al. (2025) Euclid Collaboration: Cropper, M., Al-Bahlawan, A., Amiaux, J., et al. 2025, A&A, 697, A2
- Euclid Collaboration: Jahnke et al. (2025) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2025, A&A, 697, A3
- Euclid Collaboration: Mellier et al. (2025) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2025, A&A, 697, A1
- Euclid Collaboration: Scaramella et al. (2022) Euclid Collaboration: Scaramella, R., Amiaux, J., Mellier, Y., et al. 2022, A&A, 662, A112
- Euclid Collaboration: Schirmer et al. (2022) Euclid Collaboration: Schirmer, M., Jahnke, K., Seidel, G., et al. 2022, A&A, 662, A92
- Grupp et al. (2012) Grupp, F., Prieto, E., Geis, N., et al. 2012, in Space Telescopes and Instrumentation 2012: Optical, Infrared, and Millimeter Wave, Vol. 8442 (SPIE), 358–368
- Jiang et al. (2016) Jiang, J., Lei, L., & Guangjun, Z. 2016, Optical Engineering, 55, 063101, publisher: SPIE
- Korsch (1977) Korsch, D. 1977, Applied Optics, 16, 2074
- Kramida et al. (2022) Kramida, A., Ralchenko, Y., Reader, J., & Team, N. A. 2022, National Institute of Standards and Technology, Gaithersburg, MD., version 5.10 [Online]
- Laureijs et al. (2011) Laureijs, R., Amiaux, J., Arduini, S., et al. 2011, arXiv:1110.3193
- Le Graët et al. (2022) Le Graët, J., Secroun, A., Barbier, R., et al. 2022, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 12191, X-Ray, Optical, and Infrared Detectors for Astronomy X, ed. A. D. Holland & J. Beletic, 121911M
- Lucy (1974) Lucy, L. B. 1974, AJ, 79, 745
- Maciaszek et al. (2022) Maciaszek, T., Ealet, A., Gillard, W., et al. 2022, in Space telescopes and instrumentation 2022: Optical, infrared, and millimeter wave, ed. L. E. Coyle, S. Matsuura, & M. D. Perrin, Vol. 12180 (SPIE), 121801K
- Maciaszek et al. (2016) Maciaszek, T., Ealet, A., Jahnke, K., et al. 2016, in Space Telescopes and Instrumentation 2016: Optical, Infrared, and Millimeter Wave, ed. H. A. MacEwen, G. G. Fazio, M. Lystrup, N. Batalha, N. Siegler, & E. C. Tong, Vol. 9904, International Society for Optics and Photonics (SPIE), 99040T
- Moffat (1969) Moffat, A. F. J. 1969, A&A, 3, 455
- Moore et al. (2004) Moore, K. E., Nicholson, M. G., & Warwick, J. R. 2004, in Optical Design and Engineering, ed. L. Mazuray, P. J. Rogers, & R. Wartmann, Vol. 5249 (SPIE), 400–407
- Refregier et al. (2010) Refregier, A., Amara, A., Kitching, T. D., et al. 2010, Euclid Imaging Consortium Science Book, arXiv:1001.0061 [astro-ph]
- Richardson (1972) Richardson, W. H. 1972, J. Opt. Soc. Am., JOSA, 62, 55
- Serre et al. (2010) Serre, D., Villeneuve, E., Carfantan, H., et al. 2010, in Adaptive Optics Systems II, Vol. 7736 (SPIE), 1498–1509
Appendix A On ground calibration coefficients
Table 2 presents the averaged calibration coefficients used to model the spectral dispersion of the NISP grism. The table provides the Chebyshev polynomial coefficients directly derived from a fit of Eq. 3 to the ground dataset for all grisms but RGS270 and averaged over the NISP focal plane surface. For convenience, the table provides as well the equivalent averaged coefficients for regular polynomials. Those were obtained through the analytical transformation from the Chebyshev basis into the regular polynomial basis.
Table 3 to 9 give the calibration coefficients for Eq. 3, used to model special dependency of the coefficient needed in Eq. 2.
Grism | Coef. | Chebychev polynomial coefficients | Regular polynomial coefficient | ||||||
BGS000 | |||||||||
RGS000-4 | |||||||||
RGS000 | |||||||||
RGS000+4 | |||||||||
RGS180-4 | |||||||||
RGS180 | |||||||||
RGS180+4 | |||||||||